OpenStudy (richyw):

Write an equation for the plane tangent at the point \(P(2,1,1)\) to the surface with the equation \(x^3+y^3+z^3)=5xyz\) So I have implicitly differentiated this wrt x and y. And obtained \(z_x=1\) and \(z_y=-1\) at the given point. I'm just a little confused how to use that to get the plane equation!

OpenStudy (anonymous):

well u have z=Ax+By+C for your plane am i right?

OpenStudy (anonymous):

the equations \(\large z_{x}=1 \) and \(\large z_{y}=-1 \) also must satisfy the plane

OpenStudy (richyw):

hmm. well usually I would do \(z-z_0=f_x(x_0,y_0)(x-x_0)+f_y(x_0,y_0)(y-y_0)\) but with this implicit differentiation I am lost

OpenStudy (richyw):

so I think I see that that would be \(z-z_0=(x-2)-(y-1)\). Is that correct so far? If so would the wohle thing just be\[z-1=x-y-1\]\[x-z-y=0\]

OpenStudy (anonymous):

thats right